C=100+0.80Y,what is the savings function in the simple Keynesian model?
1. S=-100+0.20Y
2. S= 100+0.20Y
3. 0.20
4.5
1
Expert's answer
2020-04-20T09:55:00-0400

Saving is the rest of income after consumption. So we cam find saving curve:

"S=Y-C=Y-100-0.8Y=0.2Y-100."

Saving function:

"S = 0.2Y-100."

So the right answer is 1.